A large pump can fill a 54,000 gallon reservoir in 8 hours. Working together, the largepump and a smaller one can fill the reservoir in 6 hours. How long would it take the smaller pump to fill the reservoir by itself?
: A large pump can fill a 54,000 gallon reservoir in 8 hours. Working together, the largepump and a smaller one can fill the reservoir in 6 hours. How long would it take the smaller pump to fill the reservoir by itself?

Answer by jojo14344(809)

Remember:
LargePump=highlight(LP)=54000gal/8hrs ----> condition 1
SmallPump=highlight(SP)=unknown
But, LP+SP=54000gal/6hrs ---------------> condition 2
Substitute condtn. 1 in condtn. 2:
(54000/8)+SP=54000/6
SP=(54000/6)-(54000/8). Take LCD,
SP=(432000-324000)/48=(108000/48) Divide both numerator/denominator by 2:
SP=cross(108000)54000/cross(48)24
SP=54000gal/24hrs
Thank you,
Jojo